A boat travelled at a certain speed to a jetty 30 km away, then reduced speed by 4km/hr and returned to its starting
point.
If the boat had travelled at a steady and constant speed of 8km/hr for the entire trip, the total time would have been
half an hour less.
Determine the initial speed of the boat.

Answer by jorel555(1290)

If the boat had a constant speed of 8 km/h over 60 kilometers, it would have taken 60/8 hours, or 7.5 hours. A half hour more would make it 8 hours. Let r be the boat's returning rate. Then:
30/(r+4)+30/r=8
30r+30(r+4)=8(rē+4r)
60r+120=8rē+32r
8rē-28r-120=0
2rē-7r-30=0
(2r+5)(r-6)=0
r=-5/2 or 6
Throwing out the negative result, we get a returning rate of 6 km/h, which makes the starting rate 10 km/h.